Day shift summary

Nickolas Solomey (solomey@hep.uchicago.edu)
Fri, 24 Sep 1999 15:26:53 -0500

Shift Summary for day shift, 24/09/99

On Shift: N. Solomy, E. Monnier

Hours of beam delivered at all intensities: 3.75
Hours of beam delivered at nominal intensity: 3.75
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Tape changing, and no beam from accelerator.

========================================================================
Runs ended this shift: 14703 long run 14.8 M events,
but very little data in any subsequent run and then the
power substation to MI went out.

Tapes brought to FCC this shift: UP(A,B,C)022-024 and UPS008 (Thanks,
Hogan for taking them to the Feynman center over lunch!! a kind sole
on KTeV)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift: Cleaned the refrigerator, almost everything thrown
away, but in the shift's opinion it was not ediable anyway so no loss.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14703 7:05 New shift arrives a tad late.
9:10 Ed called, he changed the prescale for ke3 accepts
from 12 to 8 so there are more of them for
calibration, it will take effect for next run.
9:15 MCR called said that autotune is runnig from
KTeV and they want to run it from the control room
so they asked us to stop ours.
309 KTEVANA errors so we ran KFREND
310 Tapes full, run auto terminate.
14704 10:00 Run started, but no beam now.
run died due to pipeline errors and since there
was no beam we terminated.
10:30 Bob T. called MCR to ask for 1e13 when beam
returns.
14706 Again pipeline problems so we scraped it.
14707 11:25 Run started, still no beam, power supply problems.
11:43 Ashkan called, since no beam he asked up to stop
run and update constants.
14708 11:56 Run started beam is now back we missed 2 spills.
4 Beam gone
5 12:02 Beam back
10 12:08 Beam seems to have been stable for last 3 spills
at 8e12.
15 Beam gone.
20 Beam back, 5e12.
27 9e12, they are trying to get our intensity up.
28 8.8e12
32 12:40 Power outage at Kautz Road substation, beam lost.
13:05 Restarted SDAQ with fresh start because all
windows except HV monitoring went dead.
13:30 Run stopped, Theo suggested, to resolve the sdaq
problem communication problem with the DA that
was caused by the power outage.
14709 14:20 Run starting signals from accelertor are back but
no beam expected for a while still.
14:45 Problem with DA, strobe signals from accelertor
back but DA won't even run laser triggers,
Paged Bob T.
15:00 Tried DRT_fresh_start but have pipeline crate
8 error, Paged Bob T. but Hogan answered the
Page, he walked us threw a crate 8 boot, but
still no problem.
15:05 Bob T. phone back while Hogan was working on
problem by phone with us, he is coming over
claiming he knows what the problem is, still no
sign of beam.
15:15 Talked with MCR, they lost a lot of Helium but
not enough to warm up the accelertor. They will
be taking at least 4 hours to recover.
15:30 End of shift, next shift has to get DA going.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey